# Unit-generated orders of real quadratic fields I. Class number bounds

### v2

@@ Abstract.
Unit-generated orders of a quadratic field are orders of the form $\mathcal{O}=\mathbb{Z}[\varepsilon]$, where $\varepsilon$ is a unit in the quadratic field. If the order $\mathcal{O}$ is a maximal order of a real quadratic field, then the quadratic number field is necessarily of a restricted form, being of narrow Richaud–Degert type. However, every real quadratic field contains infinitely many distinct unit-generated orders. They are parametrized as $\mathcal{O}=\mathcal{O}_{n}^{\pm}$ having quadratic discriminants $\Delta(\mathcal{O})=\Delta_{n}^{+}=n^{2}-4$ (for $n\geq 3$) and $\Delta(\mathcal{O})=\Delta_{n}^{-}=n^{2}+4$ (for $n\geq 1$). We show the (wide or narrow) class numbers of unit-generated orders satisfy $\log\left|{\rm Cl}(\mathcal{O})\right|\sim\log\frac{1}{2}\left|\Delta(\mathcal{O})\right|$ as $\left|\Delta(\mathcal{O})\right|\to\infty$, using a result of L.-K. Hua. We deduce that there are finitely many unit-generated quadratic orders of class number one and finitely many unit-generated quadratic orders whose class group is $2$-torsion. We classify all unit-generated real quadratic maximal orders having class number one. We provide numerical lists of quadratic unit-generated orders whose class groups are $2$-torsion for $\Delta\leq 10^{10}$, for both wide and narrow class groups, which are conjecturally complete.
@@ 1. Introduction
    A complete classification of unit-generated maximal orders having class number one.
@@
    A list of unit-generated orders (maximal and non-maximal) having class number one, which is complete for $\Delta<10^{10}$ and conjecturally complete overall.
  4. (4)
    A list of unit-generated orders (maximal and non-maximal) whose wide class group is $2$-torsion, which is complete for $\Delta<10^{10}$. It identifies all such orders having one class per genus in this range. It is conjecturally complete for both problems.
@@ 1.1. Results
All of the orders in these parameter ranges are distinct, with one exception: $\mathcal{O}_{\Delta^{+}_{3}}=\mathcal{O}_{\Delta^{-}_{1}}$ each give the maximal order of $\mathbb{Q}(\sqrt{5})$. For the remaining parameter values in the range $n\geq 0$, $\Delta_{2}^{+}=0$ and $\Delta_{0}^{-}=4$ are not discriminants of quadratic orders.
@@ 1.1.1. Class number bounds
It is known that there exist infinitely many (non-maximal) real quadratic orders of class number one. The first examples were given by Dirichlet [18] in 1855 in the context of binary quadratic forms. For example, $\Delta=5^{2k+1}$ for $k\geq 0$ are discriminants of orders in $\mathbb{Q}(\sqrt{5})$, which all have $h_{\Delta}=1$. (See [25, Theorem 5.9.12] and [50, Section 9, Aufgabe 5]).
@@
1.2 is an immediate consequence of a theorem of Hua [27, Theorem 12.15.4], which gives an extension to quadratic orders of Siegel’s original version of the Brauer–Siegel theorem. The error term in this result is ineffective.
3.3(2) below shows that if one restricts to families of unit-generated orders having bounded associated fundamental discriminants $(\Delta_{n}^{\pm})_{0}\leq N$, then there is an effective remainder term $O_{\!N}(\log\log n)$. Bounded families arise in the context of SIC-POVMs (discussed in Section 1.2), for example, in the Fibonacci–Lucas family of SICs studied in [23].
@@ 1.1.2. Class number one
From Theorem 1.2 we have:
Corollary 1.3.
There are finitely many unit-generated real quadratic orders having class number one.
Proof.
If there were infinitely many, then for that set of $\Delta^{\pm}_{n}$ we immediately contradict (1.1) or (1.2). ∎
1.3 leads to the following problem.
Problem 1.4.
Determine all unit-generated quadratic orders $\mathcal{O}_{\Delta}$ having $h_{\Delta}=1$.
Classifying the set of discriminants of unit-generated real quadratic orders having class number one is analogous to the famous Gauss problem of classifying all discriminants $\Delta<0$ having class number one, an analogy we discuss further in Section 5.
This problem has been well-studied for maximal orders. The next result gives a complete classification for unit-generated orders that are maximal orders.
@@
The maximal orders of discriminant $\Delta_{n}^{+}=n^{2}-4$ having class number one are those for $n\in\{0,1,3,4,5,9,21\}$, having $\Delta\in\{-4,-3,5,12,21,77,437\}$. The maximal orders of discriminant $\Delta_{n}^{-}=n^{2}+4$ having class number one are those for $n\in\{1,2,3,5,7,13,17\}$, having $\Delta\in\{5,8,13,29,53,173,293\}$.
1.5 is proved in Section 4.1 by combining four results. The first is a major result of Biró [9] handling the case $n^{2}+4$ with $n$ odd. The second is a a result of Byeon, Kim, and Lee [10] handling the case $n^{2}-4$ with $n$ odd, using a modification of Biró’s method. This paper completes the classification in the remaining two cases, $n^{2}+4$ with $n$ even (Theorem 4.2) and $n^{2}-4$ with $n$ even (Theorem 4.5), using genus theory.
There are eleven real quadratic fields with unit-generated maximal order having class number one, compared with nine imaginary quadratic fields having class number one. In Section 4.2, we provide an extended table (Table 2) of all unit-generated orders of class number one, including non-maximal orders for discriminants $\Delta<10^{10}$.
@@ 1.1.3. One class per genus and $2$-torsion class group
The genus theory of Gauss extends to all (real and imaginary) quadratic orders. The principal genus theorem of Gauss asserts that the (narrow) ideal classes in the principal genus are the squares of all ideal classes. Two ideal classes $\mathfrak{a}]$ and $\mathfrak{b}]$ in the narrow ideal class group $\operatorname{Cl}^{+}(\mathcal{O})$ correspond under Gauss composition to binary quadratic form classes of the same genus if and only if $\mathfrak{a}\mathfrak{b}^{-1}]\in\operatorname{Cl}^{+}(\mathcal{O})^{2}$. The genus group is $\operatorname{Cl}^{+}(\mathcal{O})/\operatorname{Cl}^{+}(\mathcal{O})^{2}$, which is isomorphic to the $2$-torsion subgroup $\operatorname{Cl}^{+}(\mathcal{O})[2]$. (See [25, Sections 5.6 and 6.5].)
In Section 3 we establish the following result, which strengthens 1.3.
@@
The following corollary follows immediately.
@@ 1.3. Prior work
For background on the theory of real quadratic orders, see the books of Halter-Koch [25, Chapter 5] and Mollin [40, Chapters 3 and 5].
@@
In 1996, Mollin [40, pp. 77–78] introduced the terminology narrow Richaud–Degert type for the special case of $D=m^{2}+r$ with $r\in\{\pm 1,\pm 4\}$, where $D$ is required to be squarefree. In 1998 Mollin [41] considered $\Delta^{\pm}$ as parametric families of quadratic orders, allowing the non-maximal order case. He gave a sufficient condition for class number one.
Mollin [40, pp. 77–78] also introduced the terminology extended Richaud–Degert type, abbreviated ERD-type, for
@@
removing Hasse’s restriction m<r<m-m<r<m; see also [42]. All fields of ERD-type contain unit-generated orders of small conductor (that is, small index in the maximal order). This fact can be deduced from [40, Theorem 3.2.1], which presents the continued fractions of some $\omega$ with $\mathcal{O}_{K}=\mathbb{Z}[\omega]$ in all possible cases in which $K$ is of ERD-type. In this paper, we give the largest unit-generated order in fields of narrow Richaud–Degert type in 2.5 and Table 1. In these fields, the largest unit-generated order has conductor either $1$ or $2$.
General criteria for real quadratic orders to have class number one are given recently by Kawamoto and Kishi [29, Theorem 2.6], [30, Theorem 3.5].
@@ 1.4. Contents of paper
Section 3 proves asymptotic results on class numbers $h_{\Delta}=\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|$ for $\Delta=n\pm 4$. It proves finiteness of such $n$ for which $\mathcal{O}_{\Delta}$ has one class per genus.
Section 4 treats the problem of finding all such $\Delta=n^{2}\pm 4$ having $h_{\Delta}=1$. It gives a complete answer for the case of maximal orders, where $\Delta$ is a fundamental discriminant. This section also gives a list of known unit-generated orders for which the class number is $1$; this list is conjecturally complete and known to be complete for discriminants below $10^{10}$.
Section 5 discusses the problem of finding all $\Delta=n^{2}\pm 4$ such that $\mathcal{O}_{\Delta}$ has one class per genus. It gives lists of known unit-generated order discriminants for which the (wide) class group is $2$-torsion; these lists are conjecturally complete and are known to be complete for discriminants below $10^{10}$. It also determines the subset of those having one class per genus (those whose narrow class group is $2$-torsion).
@@ 2. Basic properties of unit-generated orders
$\Delta_{0}=\begin{cases}D_{0}&\mbox{if}\quad D_{0}\equiv 1\,\,\left(\operatorname{mod}\,4\right)\\ 4D_{0}&\mbox{if}\quad D_{0}\equiv 2,3\,\,\left(\operatorname{mod}\,4\right).\end{cases}$ (2.2)
Following Mollin [40, p. 4], we term $D_{0}$ the fundamental radicand of $\Delta_{0}$ (or of $\mathbb{Q}(\sqrt{\Delta_{0}})$).
@@ 2.2. Relation between narrow Richaud–Degert fields and unit-generated orders
Proposition 2.5.
  • (1)
  • (2)
Case (2) occurs if and only if $D_{0}=m^{2}+1$ and $m$ is even. Conversely, if $K$ is a real quadratic field whose maximal order is unit-generated, then $K$ is of narrow Richaud–Degert type.
Proof.
The proposition follows by checking the five cases illustrated in Table 1. Inspection of the conditions $D_{0}\equiv 1,2,3\,\,\left(\operatorname{mod}\,4\right)$ and $\Delta_{0}\equiv 0,1\,\,\left(\operatorname{mod}\,4\right)$ shows that the rows of this table exhaust both all $D_{0}$ of narrow Richaud–Degert type and all $\Delta_{0}$ of maximal unit-generated orders. ∎
@@ 2.3. Class groups of orders
The (wide) class group (also called the Picard group or ring class group) of an order $\mathcal{O}$ is
@@
Since a quadratic order is uniquely determined by its discriminant $\Delta$, we will sometimes write $\operatorname{Cl}(\Delta)$ as an abbreviation for $\operatorname{Cl}(\mathcal{O}_{\Delta})$.
The wide class group quotients by all principal fractional ideals, while the narrow class group $\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})$, in the real quadratic case, quotients by the subgroup of all principal ideals having a totally positive generator, equivalently, having a generator of positive norm. Thus
@@
We have $\frac{\left|\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})\right|}{\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|}=1$ or $2$, the former occurring exactly when there is a unit of norm 1-1 in the order $\mathcal{O}_{\Delta}$.
Nonmaximal orders always contain non-invertible integral ideals and noninvertible fractional ideals. If one instead considers all nonzero fractional ideals, quotienting by nonzero principal ideals (which are always invertible), one obtains a monoid (that is, a semigroup with identity), the class monoid $\operatorname{Clm}(\mathcal{O})$. The class monoid is a group when $\mathcal{O}=\mathcal{O}_{K}$, and is not a group otherwise. We consider class monoids for unit-generated orders in a sequel [33].
@@ 3.2. Growth of class numbers of unit-generated orders
We write unit-generated order $\Delta_{n}^{\pm}=f_{\Delta}^{2}\Delta_{0}$, where $\Delta_{0}$ is fundamental. We set $f=f_{\Delta}$ and use the formula for the class number of an order given in Neukirch [44, Theorem 12.12]:
@@
see [46, Theorem 15] and [6, Theorem 8.8.7]. Applying this estimate (on both sides) yields
@@
3.3. Finiteness of unit-generated orders with one class per genus
We recall facts about the class group $\operatorname{Cl}(\mathcal{O}_{\Delta})$ of an order $\mathcal{O}_{\Delta}$ and the genus theory of Gauss for general binary quadratic forms, described in [25, Chapter 6.5].
  1. (1)
  2. (2)
  3. (3)
Then its order is given by [25, Theorem 6.5.2], which we now restate.
Proposition 3.4.
If $\Delta$ is a quadratic discriminant, then
We observe that $\mu(\Delta)-1\leq\omega(\Delta)$, where $\omega(m)$ counts the number of distinct prime divisors of $m$ (without multiplicity). It is well known that $\omega(m)=O\!\left(\frac{\log m}{\log\log m}\right)$ for $m\geq 2$ [43, Theorem 2.10], so we conclude, for $\Delta\geq 20$,
$\mu(\Delta)-1=O\!\left(\frac{\log\Delta}{\log\log\Delta}\right).$ (3.29)
The constant in the $O$-notation is effectively computable.
Proof of 1.6.
We first consider general quadratic discriminants $\Delta$ (not necessarily unit-generated) such that $\operatorname{Cl}(\mathcal{O}_{\Delta})$ is $2$-torsion. Then by 3.4 and (3.29),
$\log\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|\leq\log\left|\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})\right|=\log\left|\mathfrak{F}_{\Delta}[2]\right|=\log 2^{\mu(\Delta)-1}=O\!\left(\frac{\log\Delta}{\log\log\Delta}\right).$ (3.30)
From (3.30) it follows that, for any infinite family $\mathcal{F}$ of such orders in which $\operatorname{Cl}(\mathcal{O}_{\Delta})$ is $2$-torsion, we necessarily have
Such infinite families do exist since, as noted earlier, there exist infinitely many real quadratic orders of class number $1$.

### v3

@@ Abstract.
Unit-generated orders of a quadratic field are orders of the form $\mathcal{O}=\mathbb{Z}[\varepsilon]$, where $\varepsilon$ is a unit in the quadratic field. If the order $\mathcal{O}$ is a maximal order of a real quadratic field, then the quadratic number field is necessarily of a restricted form, being of narrow Richaud–Degert type. However, every real quadratic field contains infinitely many distinct unit-generated orders. They are parametrized as $\mathcal{O}=\mathcal{O}_{n}^{\pm}$ having quadratic discriminants $\Delta(\mathcal{O})=\Delta_{n}^{+}=n^{2}-4$ (for $n\geq 3$) and $\Delta(\mathcal{O})=\Delta_{n}^{-}=n^{2}+4$ (for $n\geq 1$). We show the (wide or narrow) class numbers of unit-generated orders satisfy $\log\left|{\rm Cl}(\mathcal{O})\right|\sim\log\frac{1}{2}\left|\Delta(\mathcal{O})\right|$ as $\left|\Delta(\mathcal{O})\right|\to\infty$, using a result of L.-K. Hua. We deduce that there are finitely many unit-generated quadratic orders of class number one and finitely many unit-generated quadratic orders whose class group is $2$-torsion. We classify all unit-generated real quadratic orders having class number one. We provide numerical lists of quadratic unit-generated orders whose class groups are $2$-torsion for $\Delta\leq 10^{10}$, for both wide and narrow class groups. These lists are conjecturally complete for all $\Delta$.
@@ 1. Introduction
    A complete classification of unit-generated quadratic orders having class number one.
@@
    An (ineffective) proof of finiteness of the number of unit-generated orders whose wide class group is $2$-torsion. This set includes all unit-generated orders having one class per genus. We tabulate a list of unit-generated quadratic orders whose wide class group is $2$-torsion, complete for $\Delta<10^{10}$, and identify all such orders having one class per genus for $\Delta<10^{10}$. The list is conjecturally complete for both problems.
@@ 1.1. Results
All of the orders in these parameter ranges are distinct, with one exception: The orders $\mathcal{O}_{\Delta^{+}_{3}}=\mathcal{O}_{\Delta^{-}_{1}}$ are both the maximal order of $\mathbb{Q}(\sqrt{5})$. For the remaining parameter values in the range $n\geq 0$, $\Delta_{2}^{+}=0$ and $\Delta_{0}^{-}=4$ are not discriminants of quadratic orders.
@@ 1.1.1. Class number bounds
It is known that there exist infinitely many (non-maximal) real quadratic orders of class number one. The first examples were given by Dirichlet [18] in 1855 in the context of binary quadratic forms. For example, $\Delta=5^{2k+1}$ for $k\geq 0$ are discriminants of orders in $\mathbb{Q}(\sqrt{5})$, which all have $h_{\Delta}=1$. (See [25, Theorem 5.9.12] and [49, Section 9, Aufgabe 5]).
@@
This result is proved as 3.3(1). It is an immediate consequence of a theorem of Hua [27, Theorem 12.15.4], which gives an extension to quadratic orders of Siegel’s original version of the Brauer–Siegel theorem. The error term in this result is ineffective.
3.3(2) shows that if one restricts to families of unit-generated orders having bounded associated fundamental discriminants $(\Delta_{n}^{\pm})_{0}\leq N$, then there is an effective remainder term $O_{\!N}(\log\log n)$. Bounded families arise in the context of SIC-POVMs (discussed in Section 1.2), for example, in the Fibonacci–Lucas family of SIC-POVMs studied in [23].
@@ 1.1.2. Class number one
We give a complete classification of all unit-generated quadratic orders having class number one.
Classifying the set of discriminants of unit-generated real quadratic orders having class number one is analogous to the famous Gauss problem of classifying all discriminants $\Delta<0$ having class number one. The imaginary quadratic discriminant problem for maximal orders was solved by the Heegner–Stark–Baker Theorem. Stark [47] notes that the non-maximal order case for imaginary quadratic fields is solved as a consequence of the maximal order case, giving the four discriminants $\{-12,-16,-27,-28\}$, using the fact that non-maximal orders can have class number one only if the maximal order also has class number one.
Our first result gives a complete classification for unit-generated quadratic orders that are maximal orders.
@@
The maximal quadratic orders of discriminant $\Delta_{n}^{+}=n^{2}-4$ having class number one are those for $n\in\{0,1,3,4,5,9,21\}$, having $\Delta\in\{-4,-3,5,12,21,77,437\}$. The maximal quadratic orders of discriminant $\Delta_{n}^{-}=n^{2}+4$ having class number one are those for $n\in\{1,2,3,5,7,13,17\}$, having $\Delta\in\{5,8,13,29,53,173,293\}$.
1.3 is proved in Section 4.1 by combining four results. The first is a major result of Biró [9] handling the case $n^{2}+4$ with $n$ odd. The second is a a result of Byeon, Kim, and Lee [10] handling the case $n^{2}-4$ with $n$ odd, using a modification of Biró’s method. This paper completes the classification in the remaining two cases, $n^{2}+4$ with $n$ even (4.2) and $n^{2}-4$ with $n$ even (4.5), using genus theory.
The next result gives a complete classification for unit-generated orders that are non-maximal orders.
Theorem 1.4.
The non-maximal quadratic orders of discriminant $\Delta_{n}^{+}=n^{2}-4$ having class number one are those for $n\in\{6,7,11\}$, having $\Delta\in\{32,45,117\}$. The non-maximal quadratic orders of discriminant $\Delta_{n}^{-}=n^{2}+4$ having class number one are those for $n\in\{4,8,11\}$, having $\Delta\in\{20,68,125\}$.
The proof of 1.4 builds on 1.3. The idea of the proof is similar to the imaginary quadratic order case, in which all the over-orders of a non-maximal imaginary quadratic order having class number one must also be unit-generated orders. In the real quadratic unit-generated case, however, there is an additional exceptional case; see Section 4.2. The exceptional case is covered by a major result of Biró [8] solving Chowla’s conjecture for maximal orders with $\Delta=4n^{2}+1$.
@@ 1.1.3. One class per genus and $2$-torsion class group
The genus theory of Gauss extends to all (real and imaginary) quadratic orders. The principal genus theorem of Gauss asserts that the (narrow) ideal classes in the principal genus are the squares of all ideal classes. Two ideal classes $\mathfrak{a}]$ and $\mathfrak{b}]$ in the narrow ideal class group $\operatorname{Cl}^{+}(\mathcal{O})$ (defined in Section 2.3) correspond under Gauss composition to binary quadratic form classes of the same genus if and only if $\mathfrak{a}\mathfrak{b}^{-1}]\in\operatorname{Cl}^{+}(\mathcal{O})^{2}$. The genus group is $\operatorname{Cl}^{+}(\mathcal{O})/\operatorname{Cl}^{+}(\mathcal{O})^{2}$, which is isomorphic to the $2$-torsion subgroup $\operatorname{Cl}^{+}(\mathcal{O})[2]$. (See [25, Sections 5.6 and 6.5].)
In Section 5.2, we establish the following result regarding the wide class group and its $2$-torsion subgroup.
@@
This result is ineffective; it depends on the ineffective result of Hua [27, Theorem 12.15.4]. It has the following corollary for the one class per genus problem.
@@ 1.3. Prior work
For background on the theory of real quadratic orders, see the books of Halter-Koch [25, Chapter 5] and Mollin [39, Chapters 3 and 5].
@@
In 1996, Mollin [39, pp. 77–78] introduced the terminology narrow Richaud–Degert type for the special case of $D=m^{2}+r$ with $r\in\{\pm 1,\pm 4\}$, where $D$ is required to be squarefree. In 1998 Mollin [40] considered $\Delta^{\pm}$ as parametric families of quadratic orders, allowing the non-maximal order case. He gave a sufficient condition for class number one.
Mollin [39, pp. 77–78] also introduced the terminology extended Richaud–Degert type, abbreviated ERD-type, for
@@
removing Hasse’s restriction m<r<m-m<r<m; see also [41]. All fields of ERD-type contain unit-generated orders of small conductor (that is, small index in the maximal order). This fact can be deduced from [39, Theorem 3.2.1], which presents the continued fractions of some $\omega$ with $\mathcal{O}_{K}=\mathbb{Z}[\omega]$ in all possible cases in which $K$ is of ERD-type. In this paper, we give the largest unit-generated order in fields of narrow Richaud–Degert type in 2.5 and Table 1. In these fields, the largest unit-generated order has conductor either $1$ or $2$.
General criteria for real quadratic orders to have class number one were given recently by Kawamoto and Kishi [29, Theorem 2.6], [30, Theorem 3.5]. Caeiro and Darmon [11, Main Theorem, Theorem 25, and Theorem 28] present conditional results classifying all unit-generated orders of class number one modulo their Conjecture 7. Their Conjecture 7 asserts that the image of a real multiplication (RM) point under a certain rigid analytic cocycle may be used to define a global point on an elliptic curve defined over a ring class field, and it also asserts the truth of a Gross–Zagier formula in this context.
@@ 1.4. Contents of paper
Section 3 proves asymptotic results on class numbers $h_{\Delta}=\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|$ for $\Delta=n\pm 4$ as $n\to\infty$. It proves effective bounds for class numbers restricted to families of such orders that have bounded associated fundamental discriminants.
Section 4 classifies all discriminants $\Delta=n^{2}\pm 4$ of orders having $h_{\Delta}=1$.
Section 5 proves (ineffectively) the finiteness of the set of $\Delta=n^{2}\pm 4$ such that $\operatorname{Cl}(\mathcal{O}_{\Delta})$ is $2$-torsion. It thus proves finiteness of the set of such $\Delta$ with one class per genus. It gives lists of known unit-generated order discriminants for which the (wide) class group is $2$-torsion; these lists are conjecturally complete and are known to be complete for discriminants below $10^{10}$. It also determines the sublist of those discriminants below $10^{10}$ having one class per genus (those whose narrow class group is $2$-torsion).
@@ 2. Basic properties of unit-generated orders
$\Delta_{0}=\begin{cases}D_{0}&\mbox{if}\quad D_{0}\equiv 1\,\,\left(\operatorname{mod}\,4\right),\\ 4D_{0}&\mbox{if}\quad D_{0}\equiv 2,3\,\,\left(\operatorname{mod}\,4\right).\end{cases}$ (2.2)
Following Mollin [39, p. 4], we term $D_{0}$ the fundamental radicand of $\Delta_{0}$ (or of $\mathbb{Q}(\sqrt{\Delta_{0}})$).
@@ 2.2. Relation between narrow Richaud–Degert fields and unit-generated orders
Proposition 2.5.
  • (1)
  • (2)
Case (2) occurs if and only if $D_{0}=m^{2}+1$ and 2|m2\,|\,m. Conversely, if $K$ is a real quadratic field whose maximal order is unit-generated, then $K$ is of narrow Richaud–Degert type.
Proof.
The proposition follows by checking the five cases illustrated in Table 1. Inspection of the conditions $D_{0}\equiv 1,2,3\,\,\left(\operatorname{mod}\,4\right)$ and $\Delta_{0}\equiv 0,1\,\,\left(\operatorname{mod}\,4\right)$ shows that the rows of this table exhaust both all $D_{0}$ of narrow Richaud–Degert type and all $\Delta_{0}$ of maximal unit-generated quadratic orders. ∎
@@ 2.3. Class groups of orders
The (wide) class group (also called the Picard group or ring class group) of an order $\mathcal{O}$ in a number field $K$ is
@@
The (wide) class number $h_{\Delta}=\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|$.
Definition 2.7.
In the real quadratic order case, the narrow class group $\operatorname{Cl}^{+}(\mathcal{O})$ is the quotient group of the set of invertible fractional ideals by the subgroup of all principal ideals having a totally positive generator. That is,
@@
(Note that one can replace $\alpha$ by $\alpha$ without changing the ideal $\alpha\mathcal{O}$, so the condition $\alpha$ is totally positive” can be replaced by either the condition $\alpha$ is totally negative” or the condition $\alpha$ has positive norm.”)
We have $\frac{\left|\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})\right|}{\left|\operatorname{Cl}(\mathcal{O}_{\Delta})\right|}=1$ or $2$, the former case occurring when there is a unit of norm 1-1 in the order $\mathcal{O}_{\Delta}$. The narrow class number $h^{+}_{\Delta}=\left|\operatorname{Cl}^{+}(\mathcal{O}_{\Delta})\right|$.
Non-maximal orders always contain non-invertible integral ideals and noninvertible fractional ideals. If one instead considers all nonzero fractional ideals, quotienting by nonzero principal ideals (which are always invertible), one obtains a monoid (that is, a semigroup with identity), called the (wide) class monoid $\operatorname{Clm}(\mathcal{O})$. The class monoid is a group when $\mathcal{O}=\mathcal{O}_{K}$ and is not a group otherwise. We consider class monoids for unit-generated orders in a sequel [33].
@@ 3.2. Growth of class numbers of unit-generated orders
We write unit-generated order $\Delta_{n}^{\pm}=f_{\Delta}^{2}\Delta_{0}$, where $\Delta_{0}$ is fundamental. We set $f=f_{\Delta}$ and use the formula for the class number of an order given in Neukirch [43, Theorem 12.12]:
@@
see [45, Theorem 15] and [6, Theorem 8.8.7]. Applying this estimate (on both sides) yields